Current Activities

  • The draft report related to the storm sewer leakage testing in the Golden Locust Drive area has been received and is currently being reviewed by staff;
  • Flushing of the FDC system is on-going and is approximately 50% complete. Operations staff noted a difficulty with accessing a portion of the FDC south of Doug Leavens Blvd. (approximately 1000 m) and options are being explored to provide better access to reach this area;
  • Staff are coordinating a topographical survey of Black Walnut Trail down to Derry Rd. and will also be contacting certain homeowners with respect to confirming basement elevations;
  • Lisgar Pump Protocol was initiated yesterday (Aug 17th). The pump posted at Black Walnut Trail and Derry Road was activated at 5:45pm and deactivated at approximately 6:15pm. Pipe flow was estimated to be at 70% capacity at this location. There were no issues at the other two pump locations.
  • Our consultant is analyzing the results of the storm sewer system leakage testing and will provide recommendations to staff on next steps for lining works;
  • Our consultant is reviewing the CCTV footage of the FDC and will provide recommendations with respect to maintenance issues, and
  • Our consultant is continuing with the design of the Trench Dewatering System at Black Walnut Trail and Cactus Gate and will evaluate options in consideration of data collected from the incident on July 14th. Staff will work with the consultant on the system design options and considerations over the next few weeks.

Next Steps 

  • Staff will continue to work with the consultant on the findings of its monitoring data from the July 14th incident to determine additional next steps for the investigation and mitigation plan.